OpenStudy (anonymous):

A children's birthday party at an indoor play center costs $50 to rent the space and $6.25 per child. Gina wants to spend no more than $100 on her son's fourth birthday party. Write an inequality for the number of children who can attend the birthday party, and solve the inequality. A) 50 + 6.25x ≤ 100; x ≤ 8 B) 50 + 6.25x ≥ 100; x ≥ 8 C) 50 + 6.25x < 100; x < 8 D) 50 + 6.25x > 100; x > 8

OpenStudy (anonymous):

I guess D?

jimthompson5910 (jim_thompson5910):

let x = # of children if x children go to the party, then you agree that it costs 6.25x dollars plus the $50 to rent the space so the total cost, call it T for now, is T = 50+6.25x

jimthompson5910 (jim_thompson5910):

"Gina wants to spend no more than $100 on her son's fourth birthday party." so $100 is the max she'll go what does that mean in terms of an inequality?

OpenStudy (anonymous):

C?

jimthompson5910 (jim_thompson5910):

why C?

OpenStudy (anonymous):

I am not good at this so I just guesses but here is what I think.. I think C because 50 and 6.25x is bigger then 100 but if I add 50 with 6.25 it would make 56.25 witch is higher or less..

jimthompson5910 (jim_thompson5910):

Is Gina allowed to go over $100?

OpenStudy (anonymous):

She wants to spend no more then 100..

jimthompson5910 (jim_thompson5910):

so $100 is the highest she can go so that means the total cost T must be either less than $100 or at $100 this means \[\Large T \le 100\]

jimthompson5910 (jim_thompson5910):

previously, we found the total cost to be T = 50 + 6.25x so, \[\Large T \le 100\] \[\Large 50+6.25x \le 100\] after you plug in the expression for T

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

No offense, but I'd appreciate if you did NOT just blindly guess

jimthompson5910 (jim_thompson5910):

I practically gave you the answer anyway

OpenStudy (anonymous):

I wasn`t guessing.. But thanks anyways...